Kutztown University Department of Mathematics

Colloquium

4:00 p.m.

October 18, 2007

Lytle Hall 218

How to Write Mathematics on the Web

Dr. Kunio Mitsuma

Kutztown University of Pennsylvania

Abstract

Many of us have our own web sites. We post syllabi, study materials, and such. Often, these are pdf, Word, and other files that force visitors to launch suitable applications in order to view them. Leaving a web browser causes discontinuity. We should strive for web sites that are contained in the html language. But, how do we make mathematics symbols and equations appear on web pages without resorting to copied-and-pasted gif images? In this talk, we will explore various ways to write mathematics on the web. MathML, html ports of TeX, Maple, and LiveMath notebooks will be discussed with their pros and cons.

As an added bonus, I will provide an easy-to-follow introduction to setting up your own web site on either mantaray (faculty) or clownfish (student) servers at KU.

3:30 p.m. refreshments served

4:00 p.m. talk begins

Kutztown Univ. of Pennsylvania Dept. of Mathematics Dr. Paul Ache, Chair